Solve for the value of x that makes the equation true.
8 + 2(–6x + 5 + 7x) –12 = 6 – (47 + 3x – 7)

Answer by ewatrrr(24785)

 
Hi
Removing Parenthesis and Gathering Like Terms
8 -12x + 10 + 14x -12 = 6 -47 -3x + 7
5x = -34 - 6
5x = -40
x = -8